# Approximating Sparse Matrices and their Functions using Matrix-vector products

### v2

@@ 3.2.2 Error estimate for BaMRAM
where $\bm{X}\in\mathbb{R}^{n\times 5}$ is a Gaussian sensing operator. As previously mentioned, making a small constant number of Gaussian measurements, here we take $5$, usually suffices in practice; see [35, 48]. We recommend computing $\delta_{RE}^{(B)}$ when the decay behaviour of $\bm{B}$ is unknown or in the context of matrix functions, when the smoothness of $f$ or the bandwidth of the underlying matrix $\bm{A}$ is unknown. Later in Section 4, we illustrate the error estimate $\delta_{RE}^{(B)}$ in a numerical experiment; see Figure 2.
@@
Here we give theoretical guarantees for BaMRAM. The analysis for BaMRAM is easier and it gives a stronger bound than SpaMRAM as the precise location of the dominant entries and the qualitative behaviour of the decay is known. The proof is motivated by Proposition 3.2 in [8].
@@
where $C>0$ and $10<\lambda<$ 1 are constants. Then the output of BaMRAM, $\widehat{\bm{B}}$, satisfies
@@
using a special case of Hölder’s inequality [34, §2.3.3]. We now bound $\left\lVert\bm{E}\right\rVert_{1}$ and $\left\lVert\bm{E}\right\rVert_{\infty}$ using (12). By the construction of $\widehat{\bm{B}}$, the $i$th row of $\bm{E}$ satisfies
@@
    When $\bm{B}=f(\bm{A})$ for a $k$-banded matrix $\bm{A}$, we can use $s=2ck+1$ for some integer constant $c>0$. This value of $s$ is motivated by the fact that if $f$ is analytic then $f(\bm{A})$ can be well approximated by a low degree polynomial $p(\bm{A})$ which has bandwidth deg$(p)k$.
@@
    In many cases, the constants $C,\lambda$ and $s_{0}$ are either unknown, implicit or too expensive to compute in applications. For matrix functions, Theorem 4 relies on the off-diagonal decay bound in [8], which contains constants that are implicit as they depend on the function $f$ and the spectral property of the matrix $\bm{A}$. Therefore, without decay bounds that include explicit constants, Theorem 4 cannot serve as a priori estimate for Algorithm 2. However, the analysis in Theorem 4 can still serve as a heuristic for when Algorithm 2 will be useful.
@@ 3.3 Extensions to Kronecker Sum
Some of the results for banded matrices and matrices with exponential off-diagonal decay can be naturally extended to Kronecker sums, which arises, for example in the finite difference discretization of the two-dimensional Laplace operator [45]. In particular, the decay bounds for functions of Hermitian matrices with Kronecker structure has been studied in [9]. Let $\bm{A}^{(1)},\bm{A}^{(2)}\in\mathbb{R}^{n\times n}$. The Kronecker sum $\mathcal{A}\in\mathbb{R}^{n^{2}\times n^{2}}$ of $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ is defined by
@@
$\bm{A}\otimes\bm{B}=\begin{bmatrix}A_{11}\bm{B}&A_{12}\bm{B}&...&A_{1n}\bm{B}\\ A_{21}\bm{B}&A_{22}\bm{B}&...&A_{2n}\bm{B}\\ \vdots&\vdots&\ddots&\vdots\\ A_{n1}\bm{B}&A_{n2}\bm{B}&...&A_{nn}\bm{B}\end{bmatrix}\in\mathbb{R}^{n^{2}\times n^{2}}$
and $\bm{I}_{n}$ is the $n\times n$ identity matrix.
Suppose $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ are both banded matrices. Then $\bm{I}_{n}\otimes\bm{A}^{(2)}\in\mathbb{R}^{n^{2}\times n^{2}}$ is a banded matrix with the same bandwidth as $\bm{A}^{(2)}$ and $\bm{A}^{(1)}\otimes\bm{I}_{n}\in\mathbb{R}^{n^{2}\times n^{2}}$ is permutation equivalent to a banded matrix with the same bandwidth as $\bm{A}^{(1)}$ because $\bm{P}(\bm{A}^{(1)}\otimes\bm{I}_{n})\bm{P}^{T}=\bm{I}_{n}\otimes\bm{A}^{(1)}$ where $\bm{P}$ is the perfect shuffle matrix [47]. With this observation, we can use the recovery technique used for banded matrices in Section 3.1 to recover $\mathcal{A}$ for banded matrices $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$. Suppose that $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ have bandwidth $k^{(1)}$ and $k^{(2)}$ respectively. Notice that to recover $\mathcal{A}$, we need to recover the off-diagonal entries of $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ to recover the off-diagonal entries of $\mathcal{A}$ and recover $A_{ii}^{(1)}+A_{jj}^{(2)}$ for all $i,j$ to recover the diagonal entries of $\mathcal{A}$. We use the following $(2+2k^{(1)}+2k^{(2)})$ matrix-vector products
@@
to recover $\mathcal{A}$. Since the top left $n\times n$ block of $\mathcal{A}$ equals $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$, which is $k^{(2)}-$banded, the first $n$ rows of the matrix-vector product with $\begin{bmatrix}\bm{I}_{n}^{\left(1+2k^{(2)}\right)}\\ \bm{0}\end{bmatrix}$ recovers $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$ exactly (Theorem 3), which in turn recovers all the off-diagonal entries of $\bm{A}^{(2)}$. Similarly, $\bm{P}^{T}\begin{bmatrix}\bm{I}_{n}^{\left(1+2k^{(1)}\right)}\\ \bm{0}\end{bmatrix}$ recovers $\bm{A}^{(1)}+A_{11}^{(2)}\bm{I}_{n}$ exactly using $\bm{P}(\bm{A}^{(1)}\otimes\bm{I}_{n})\bm{P}^{T}=\bm{I}_{n}\otimes\bm{A}^{(1)}$. Therefore we get all the off-diagonal entries of $\mathcal{A}$ from recovering $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$ and $\bm{A}^{(1)}+A_{11}^{(2)}\bm{I}_{n}$. The diagonal entries of $\mathcal{A}$ can be recovered from the diagonal entries of $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$ and $\bm{A}^{(1)}+A_{11}^{(2)}\bm{I}_{n}$ by noting
@@
For matrices with rapid off-diagonal decay, we can extend the above idea in certain cases. Focusing on the matrix exponential, we have a special relation [42, §10]
@@
The strategy for recovering banded matrices and approximating matrices with rapid off-diagonal decay can also be used as laid out above in a similar setting involving Kronecker product or sum. Examples include $f(\bm{I}\otimes\bm{A})=\bm{I}\otimes f(\bm{A})$ and the formulae for the cosine and the sine of the Kronecker sum between two matrices [42, §12].
@@ 4 Numerical Illustrations
Here we illustrate BaMRAM and SpaMRAM using numerical experiments. In all experiments, we use $20$ iterations of a polynomial Krylov method and $50$ sample points for the contour integration from method 2 of [39] to evaluate the matrix-vector products with $f(\bm{A})$. In all of the plots, the $x$-axis denotes the number of matrix-vector products $s$, i.e. $\bm{x}\mapsto f(\bm{A})\bm{x}$, and the $y$-axis denotes the relative error in the $2$-norm with respect to $f(\bm{A})$ obtained by dense arithmetic in MATLAB, i.e., expm, sqrtm and logm. Here, we plot the relative error $\|f(\bm{A})-\widehat{f(\bm{A}})\|_{2}/\|f(\bm{A})\|_{2}$ instead of the absolute error $\|f(\bm{A})-\widehat{f(\bm{A}})\|_{2}$ to provide a clearer, scale-invariant view of our numerical results. For SpaMRAM, the sparsity parameter $k$ is chosen to be $8$ times smaller than the number of matrix-vector products, i.e., $s=8k$ and the sensing operator $\bm{Y}$ is Gaussian. The experiments were conducted in MATLAB version 2021a using double precision arithmetic.
@@ 4.1 Synthetic Examples
  • Banded case: symmetric $2$-banded matrix $\bm{A}_{B}\in\mathbb{R}^{1024\times 1024}$ with its nonzero entries drawn from i.i.d. $\mathcal{N}(0,1)$. $\bm{A}_{B}$ has been rescaled to $\left\lVert\bm{A}_{B}\right\rVert_{2}=0.5$.
@@
Symmetry was enforced to ensure that the eigenvalues are real, and $\bm{A}_{B}$ and $\bm{A}_{S}$ have been rescaled to ensure that their eigenvalues are sufficiently away from $-1$, which is where $\sqrt{1+x}$ and $\log(1+x)$ have singularities. This was done to ensure that the matrix-vector products $f(\bm{A})\bm{b}$ can be computed with high accuracy with 20 Krylov steps.
Figure 2 shows the accuracy as well as how well the error estimates $\delta_{RE}^{(S)}$ and $\delta_{RE}^{(B)}$ describe the behaviour of the two algorithms. For the banded case in Figure 2a, we see that the error decays exponentially until the error is dominated by either the machine precision or the error from computing the matrix-vector products. This exponential decay rate is consistent with Theorem 4. The dotted lines indicate the error estimate $\delta_{RE}^{(B)}$. We observe that $\delta_{RE}^{(B)}$, which makes $5$ extra Gaussian measurements, estimate the error well. Therefore, we recommend computing $\delta_{RE}^{(B)}$ in practice, especially in cases when either the bandwidth of the underlying matrix or the matrix function is unknown. For the general sparse case in Figure 2b, the error decays quickly with the number of nonzero entries recovered in each row. The sparsity pattern influences the decay rate of SpaMRAM, making the behaviour of the decay rate more unexpected. This has been shown partially in Theorem 1 where the bound depends on the sparsity pattern or more specifically the cardinality of the union of the combined support locations of $\bm{I}_{n},\bm{A},...,\bm{A}^{d}$ in each row for $d$ large enough. Although the behavior of SpaMRAM is irregular, $\delta_{RE}^{(S)}$ serves as an excellent error estimate for how SpaMRAM performs.
Figure 2: The accuracy of BaMRAM and SpaMRAM using two $1024\times 1024$ synthetic symmetric matrices: a random $2$-banded matrix (left) and a random sparse matrix with nonzero density $1/1024$ (right). The dotted lines are the error estimates $\delta_{RE}^{(S)}$ and $\delta_{RE}^{(B)}$ discussed in Sections 2.1.3 and 3.2.2 respectively. The error estimates can be used to determine the behaviour of SpaMRAM and BaMRAM.
@@ 4.2 Toy Examples
In this section, we manufacture a matrix $\bm{A}$ such that $f(\bm{A})$ has a sparse structure and use our algorithms to approximate $f(\bm{A})$ using matrix-vector products with $\bm{A}$ only. We explore two functions, namely $\sqrt{\bm{A}}$ and $\log{(\bm{A})}$. For the square root function, we set $\sqrt{\bm{A}}$ equal to a sparse positive definite matrix and use our algorithms to recover $\sqrt{\bm{A}}$ via matrix-vector products with $\bm{A}=\left(\sqrt{\bm{A}}\right)^{2}$ (which here is not very sparse). We follow a similar procedure for the matrix logarithm using the formula $\mbox{exp}(\log(\bm{A}))=\bm{A}$. The aim of this experiment is to show that if $f(\bm{A})$ is sparse then we can efficiently recover $f(\bm{A})$ to high accuracy using our algorithms. We choose a sparse positive definite matrix and a banded positive definite matrix from the SuiteSparse Matrix Collection [21].
@@
In Figure 3, we see that both matrices are recovered to high accuracy using our algorithms. In Figure 3a, we observe that the matrix $\mathtt{gr\_30\_30}$ is recovered to high accuracy using about $66$ matrix-vector products, which is consistent with the fact that $\mathtt{gr\_30\_30}$ is $31$-banded, and hence has at most $63$ nonzero entries per row and column. In Figure 3b, we see that the matrix $\mathtt{Trefethen\_700}$ has been recovered to high accuracy using about $180$ matrix-vector products. This is fairly consistent with the fact that with $s=180$ matrix-vector products, we recover about $22$ nonzero entries in each row and $\mathtt{Trefethen\_700}$ has at most $19$ nonzero entries per row.

### v3

@@ 3.2.2 Error estimate for BaMRAM
where $\bm{X}\in\mathbb{R}^{n\times 5}$ is a Gaussian sensing operator. As previously mentioned, making a small constant number of Gaussian measurements, here we take $5$, usually suffices in practice; see GrattonTitley-Peloquin2018; mt20. We recommend computing $\delta_{RE}^{(B)}$ when the decay behaviour of $\bm{B}$ is unknown or in the context of matrix functions, when the smoothness of $f$ or the bandwidth of the underlying matrix $\bm{A}$ is unknown. Later in Section 4, we illustrate the error estimate $\delta_{RE}^{(B)}$ in a numerical experiment; see Figure 2.
@@
Here we give theoretical guarantees for BaMRAM. The analysis for BaMRAM is easier and it gives a stronger bound than SpaMRAM as the precise location of the dominant entries and the qualitative behaviour of the decay is known. The proof is motivated by Proposition 3.2 in Benzi2007.
@@
where $C>0$ and $0<\lambda<1$ are constants. Then the output of BaMRAM, $\widehat{\bm{B}}$, satisfies
@@
using a special case of Hölder’s inequality (matcomp, §2.3.3). We now bound $\left\lVert\bm{E}\right\rVert_{1}$ and $\left\lVert\bm{E}\right\rVert_{\infty}$ using (12). By the construction of $\widehat{\bm{B}}$, the $i$th row of $\bm{E}$ satisfies
@@
    When $\bm{B}=f(\bm{A})$ for a $k$-banded matrix $\bm{A}$, we can use $s=2ck+1$ for some integer constant $c>0$. This choice of $s$ is motivated by the fact that if $f$ is analytic then $f(\bm{A})$ can be well approximated by a low degree polynomial $p(\bm{A})$ which has bandwidth deg(p)k(p)k. It is also useful to note the dependence of the decay factor $\lambda$ on the bandwidth. If the entries of a $1$-banded matrix decay away from the main diagonal at rate $\lambda$, then for a $k$-banded matrix with comparable spectral properties, the decay rate becomes $\lambda^{1/k}$ DMS1984. This behaviour aligns with the graph distance $d(i,j)$ for general sparse matrices discussed in Section 2, and the choice $s=2ck+1$.
@@
    In many cases, the constants $C,\lambda$ and $s_{0}$ are either unknown, implicit or too expensive to compute in applications. For matrix functions, Theorem 4 relies on the off-diagonal decay bound in Benzi2007, which contains constants that are implicit as they depend on the function $f$ and the spectral property of the matrix $\bm{A}$. Therefore, without decay bounds that include explicit constants, Theorem 4 cannot serve as a priori estimate for Algorithm 2. However, the analysis in Theorem 4 can still serve as a heuristic for when Algorithm 2 will be useful.
@@ 3.3 Extensions to Kronecker Sum
Some of the results for banded matrices and matrices with exponential off-diagonal decay can be naturally extended to Kronecker sums, which arises, for example in the finite difference discretization of the two-dimensional Laplace operator LeVequeBook. In particular, decay bounds for functions of Hermitian matrices with Kronecker structure have been studied in BenziSimoncini2015. Let $\bm{A}^{(1)},\bm{A}^{(2)}\in\mathbb{R}^{n\times n}$. The Kronecker sum $\mathcal{A}\in\mathbb{R}^{n^{2}\times n^{2}}$ of $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ is defined by
@@
$\bm{A}\otimes\bm{B}=\begin{bmatrix}A_{11}\bm{B}&A_{12}\bm{B}&...&A_{1n}\bm{B}\\ A_{21}\bm{B}&A_{22}\bm{B}&...&A_{2n}\bm{B}\\ \vdots&\vdots&\ddots&\vdots\\ A_{n1}\bm{B}&A_{n2}\bm{B}&...&A_{nn}\bm{B}\end{bmatrix}\in\mathbb{R}^{n^{2}\times n^{2}}.$
Suppose $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ are both banded matrices. Then $\bm{I}_{n}\otimes\bm{A}^{(2)}\in\mathbb{R}^{n^{2}\times n^{2}}$ is a banded matrix with the same bandwidth as $\bm{A}^{(2)}$ and $\bm{A}^{(1)}\otimes\bm{I}_{n}\in\mathbb{R}^{n^{2}\times n^{2}}$ is permutation equivalent to a banded matrix with the same bandwidth as $\bm{A}^{(1)}$ because $\bm{P}(\bm{A}^{(1)}\otimes\bm{I}_{n})\bm{P}^{T}=\bm{I}_{n}\otimes\bm{A}^{(1)}$ where $\bm{P}$ is the perfect shuffle matrix VanLoan2000. With this observation, we can use the recovery technique used for banded matrices in Section 3.1 to recover $\mathcal{A}$ for banded matrices $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$. Suppose that $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ have bandwidth $k^{(1)}$ and $k^{(2)}$ respectively. Notice that to recover $\mathcal{A}$, we need to recover the off-diagonal entries of $\bm{A}^{(1)}$ and $\bm{A}^{(2)}$ to recover the off-diagonal entries of $\mathcal{A}$ and recover $A_{ii}^{(1)}+A_{jj}^{(2)}$ for all $i,j$ to recover the diagonal entries of $\mathcal{A}$. We use the following $(2+2k^{(1)}+2k^{(2)})$ matrix-vector products
@@
to recover $\mathcal{A}$. Since the top left $n\times n$ block of $\mathcal{A}$ equals $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$, which is $k^{(2)}$-banded, the first $n$ rows of the matrix-vector product with $\begin{bmatrix}\bm{I}_{n}^{\left(1+2k^{(2)}\right)}\\ \bm{0}\end{bmatrix}$ recover $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$ exactly (Proposition 3), which in turn recovers all the off-diagonal entries of $\bm{A}^{(2)}$. Similarly, $\bm{P}^{T}\begin{bmatrix}\bm{I}_{n}^{\left(1+2k^{(1)}\right)}\\ \bm{0}\end{bmatrix}$ recovers $\bm{A}^{(1)}+A_{11}^{(2)}\bm{I}_{n}$ exactly using $\bm{P}(\bm{A}^{(1)}\otimes\bm{I}_{n})\bm{P}^{T}=\bm{I}_{n}\otimes\bm{A}^{(1)}$. Therefore we get all the off-diagonal entries of $\mathcal{A}$ from recovering $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$ and $\bm{A}^{(1)}+A_{11}^{(2)}\bm{I}_{n}$. The diagonal entries of $\mathcal{A}$ can be recovered from the diagonal entries of $A_{11}^{(1)}\bm{I}_{n}+\bm{A}^{(2)}$ and $\bm{A}^{(1)}+A_{11}^{(2)}\bm{I}_{n}$ by noting
@@
For matrices with rapid off-diagonal decay, we can extend the above idea in certain cases. Focusing on the matrix exponential, we have a special relation (matfunchigham, §10)
@@
The strategy for recovering banded matrices and approximating matrices with rapid off-diagonal decay can also be used as laid out above in a similar setting involving Kronecker product or sum. Examples include $f(\bm{I}\otimes\bm{A})=\bm{I}\otimes f(\bm{A})$ and the formulae for the cosine and the sine of the Kronecker sum between two matrices (matfunchigham, §12).
@@ 4 Numerical Illustrations
Here we illustrate BaMRAM and SpaMRAM using numerical experiments. In all experiments, we use $20$ iterations of a polynomial Krylov method and $50$ sample points for the contour integration from method 2 of matveccontour to evaluate the matrix-vector products with $f(\bm{A})$. The contour integration requires the knowledge of the largest and smallest eigenvalues of $\bm{A}$. We assume that these eigenvalues are known a priori. In all of the plots, the $x$-axis denotes the number of matrix-vector products $s$, i.e. $\bm{x}\mapsto f(\bm{A})\bm{x}$. Note that this refers to the matrix-vector products with $f(\bm{A})$, not with $\bm{A}$. If computing $\bm{x}\mapsto f(\bm{A})\bm{x}$ requires $m$ matrix-vector products with $\bm{A}$, then approximating $f(\bm{A})$ in our experiments requires a total of $ms$ matrix-vector products with $\bm{A}$. The $y$-axis denotes the relative error in the $2$-norm with respect to $f(\bm{A})$ obtained by dense arithmetic in MATLAB, i.e., expm, sqrtm and logm. Here, we plot the relative error $\|f(\bm{A})-\widehat{f(\bm{A}})\|_{2}/\|f(\bm{A})\|_{2}$ instead of the absolute error $\|f(\bm{A})-\widehat{f(\bm{A}})\|_{2}$ to provide a clearer, scale-invariant view of our numerical results. For SpaMRAM, the sparsity parameter $k$ is chosen to be $8$ times smaller than the number of matrix-vector products, i.e., $s=8k$ and the sensing operator $\bm{Y}$ is Gaussian. The experiments were conducted in MATLAB version 2021a using double precision arithmetic.
@@ 4.1 Synthetic Examples
  • Banded case: symmetric $2$-banded matrix $\bm{A}_{B}\in\mathbb{R}^{1024\times 1024}$ with its nonzero entries drawn i.i.d. from $\mathcal{N}(0,1)$. $\bm{A}_{B}$ has been rescaled to $\left\lVert\bm{A}_{B}\right\rVert_{2}=0.5$.
@@
Symmetry was enforced to ensure that the eigenvalues are real, and $\bm{A}_{B}$ and $\bm{A}_{S}$ have been rescaled to ensure that their eigenvalues are sufficiently away from 1-1, which is where $\sqrt{1+x}$ and $\log(1+x)$ have singularities. This was done to ensure that the matrix-vector products $f(\bm{A})\bm{b}$ can be computed with high accuracy with 20 Krylov steps.
Figure 2 shows the accuracy as well as how well the error estimates $\delta_{RE}^{(S)}$ and $\delta_{RE}^{(B)}$ describe the behaviour of the two algorithms. For the banded case in Figure 2a, we see that the error decays exponentially until the error is dominated by either the machine precision or the error from computing the matrix-vector products. This exponential decay rate is consistent with Theorem 4. The dotted lines indicate the error estimate $\delta_{RE}^{(B)}$. We observe that $\delta_{RE}^{(B)}$, which makes $5$ extra Gaussian measurements, estimate the error well. Therefore, we recommend computing $\delta_{RE}^{(B)}$ in practice, especially in cases when either the bandwidth of the underlying matrix or the matrix function is unknown.
For the general sparse case in Figure 2b, the error decays quickly with the number of nonzero entries recovered in each row. The sparsity pattern influences the decay rate of SpaMRAM, making the behaviour of the decay rate more unpredictable and irregular. This has been shown partially in Theorem 1 where the bound depends on the sparsity pattern or more specifically the cardinality of the union of the combined support locations of $\bm{I}_{n},\bm{A},...,\bm{A}^{d}$ in each row for $d$ large enough. Although the behavior of SpaMRAM is irregular, $\delta_{RE}^{(S)}$ serves as an excellent error estimate for how SpaMRAM performs.
Finally, we note that there are regimes in which the proposed method does not convincingly amortize its computational cost. For example, in Figure 2b for the $1024\times 1024$ problem, the SpaMRAMrequires $55$ matrix–vector products with $f(\bm{A})$—corresponding to $55\times 19=1045$ matrix–vector products with $\bm{A}$—to achieve only $10^{-3}$ accuracy. In contrast, naive dense methods that explicitly form the matrix and compute a full eigenvalue decomposition achieve full accuracy using fewer matrix–vector products with $\bm{A}$. This highlights that the advantages of the proposed approach are most pronounced in large-scale settings where forming the matrix is infeasible, rather than in moderate-size problems where dense methods remain competitive. See Section 4.4 for experiments at larger scale.
Figure 2: The accuracy of BaMRAM and SpaMRAM using two $1024\times 1024$ synthetic symmetric matrices: a random 22-banded matrix (left) and a random sparse matrix with nonzero density $1/1024$ (right). The dotted lines are the error estimates $\delta_{RE}^{(S)}$ and $\delta_{RE}^{(B)}$ discussed in Sections 2.1.3 and 3.2.2 respectively. The error estimates can be used to determine the behaviour of SpaMRAM and BaMRAM.
@@ 4.2 Toy Examples
In this section, we manufacture a matrix $\bm{A}$ such that $f(\bm{A})$ has a sparse structure and use our algorithms to approximate $f(\bm{A})$ using matrix-vector products with $\bm{A}$ only. We explore two functions, namely $\sqrt{\bm{A}}$ and $\log{(\bm{A})}$. For the square root function, we set $\sqrt{\bm{A}}$ equal to a sparse positive definite matrix and use our algorithms to recover $\sqrt{\bm{A}}$ via matrix-vector products with $\bm{A}=\left(\sqrt{\bm{A}}\right)^{2}$ (which here is not very sparse). We follow a similar procedure for the matrix logarithm using the formula $\mbox{exp}(\log(\bm{A}))=\bm{A}$. The aim of this experiment is to show that if $f(\bm{A})$ is sparse then we can efficiently recover $f(\bm{A})$ to high accuracy using our algorithms. We choose a sparse positive definite matrix and a banded positive definite matrix from the SuiteSparse Matrix Collection FloridaDataset2011.
@@
In Figure 3, after an initial period of low accuracy–caused by the unrecovered entries of both matrices being $\mathcal{O}(1)$–we obtain high accuracy as the number of nonzero entries recovered per row increases sufficiently for $f(\bm{A})$, which is either sparse or banded, to be recovered exactly. In Figure 3a, the matrix $\mathtt{gr\_30\_30}$ is recovered to high accuracy using about $66$ matrix-vector products, which is consistent with the fact that $\mathtt{gr\_30\_30}$ is $31$-banded, and hence has at most $63$ nonzero entries per row and column. This corresponds to a total of $66\times 19=1254$ matrix-vector products with $\bm{A}$. In Figure 3b, the matrix $\mathtt{Trefethen\_700}$ has been recovered to high accuracy using about $180$ matrix-vector products. This is fairly consistent with the fact that with $s=180$ matrix-vector products, we recover about $22$ nonzero entries in each row and $\mathtt{Trefethen\_700}$ has at most $19$ nonzero entries per row. This corresponds to $180\times 19=3420$ matrix-vector products with $\bm{A}$.